OpenStudy (anonymous):

For which equation is x = 9 a solution? A. 3x – 7 = –34 B. x – 9 = 1 C. x/9-8=9 D. 4x + 5 = 41

OpenStudy (anonymous):

It is D.

OpenStudy (anonymous):

4x=41-5

OpenStudy (anonymous):

4x=36

OpenStudy (anonymous):

Multiply 9 by 4, and add 5, you get 41.

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

For which equation is x = 9 a solution? A. \(3x – 7 = –34 \) B. \(x – 9 = 1 \) C. \(\frac{x}{9}-8=9 \) D. \(4x + 5 = 41\) A. \(20 \neq -34\) False B. \(0 \neq 1\) False C. \(-7 \neq 9\) False D. \(41 = 41\) True How to: In an equation; say: \[4x + 5 = 41\] Input 9 for where you see \(x\). \[4(9)+5=41\] \[4(9) = 36\] \[36+5 = 41\] \[41 = 41\] Does forty one equal forty one? Yes, it does so it is true.